factory costs $460,000. You forecast that it will produce cash inflows of $150,000 in year 1, $210,000 in year 2, and $360,000 in year 3. The discount rate is 12%. a. What is the value of the factory? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

In order to calculate the cost of factory, based on the cashflows provided, we need to compute the Net Present value of the cashflows.

Mathematically,

NPV = \frac{C_0}{(1 + r)^0} + \frac{C_1}{(1 + r)^1} + \frac{C_2}{(1 + r)^2} + \frac{C_3}{(1 + r)^3}

here C0 = -460,000, C1 = 150,000, C2 = 210,000, C3 = 360,000

NPV = \frac{-460,000}{(1 + 0.12)^0} + \frac{150,000}{(1 + 0.12)^1} + \frac{210,000}{(1 + 0.12)^2} + \frac{360,000}{(1 + 0.12)^3}

NPV = $97,580.17 --> Net Value of factory today. Answer